Strickler Technology is considering changes in its working capital policies to improve its cash flow cycle. Strickler's sales last year were $2,330,000 (all on credit), and its net profit margin was 4%. Its inventory turnover was 6.5 times during the year, and its DSO was 41 days. Its annual cost of goods sold was $1,300,000. The firm had fixed assets totaling $380,000. Strickler's payables deferral period is 45 days. Assume a 365-day year. Do not round intermediate calculations.

A. Calculate Strickler's cash conversion cycle. Round your answer to two decimal places.

____ Days

B. Assuming Strickler holds negligible amounts of cash and marketable securities, calculate its total assets turnover and ROA. Round your answers to two decimal places.

Total assets turnover: _____

ROA: _____ %

C. Suppose Strickler's managers believe the annual inventory turnover can be raised to 9 times without affecting sale or profit margins. What would Strickler's cash conversion cycle, total assets turnover, and ROA have been if the inventory turnover had been 9 for the year? Round your answers to two decimal places.

Cash conversion cycle: _____ days

Total assets turnover: _____

ROA: _____ %
Answer #1

A.

Inventory Turnover Cost of Goods Sold Inventory 6.5 1,300,000 200,000 ECONOWN Days Sales Inventory Day Sales Outstanding Paya

Inventory Turnover Cost of Goods Sold Inventory 6.5 1300000 =C4/C3 Days Sales Inventory =C5/C4*365 Day Sales Outstanding 41 P

B.

12 B 6.5 Inventory Turnover Cost of Goods Sold Inventory 1,300,000 200,000 2,330,000 Total Sales Day Sales Outstanding Receiv

12 A В Inventory Turnover Cost of Goods Sold Inventory 6.5 1300000 =C4/C3 Total Sales Day Sales Outstanding Receivables 23300

C.

Inventory Turnover Cost of Goods Sold Inventory 1,300,000 144,444 2,330,000 Total Sales Day Sales Outstanding Receivables 261

Inventory Turnover Cost of Goods Sold Inventory 1300000 =C3/C2 2330000 Total Sales Day Sales Outstanding Receivables =C7/365*

Cash conversion cycle = 37 days

Total Assets Turnover = 2.96

ROA: 11.85%